OpenStudy (anonymous):

What is the solution of the equation? 4.7x + 3.8 = 13.2 need help checking my math getting 2.04

OpenStudy (anonymous):

First you would subtract 3.8 on both sides and you are left with 4.7x=9.4 Then you divide by 4.7 on both sides and get x=2.

OpenStudy (anonymous):

4.7x=13.2-3.8 4.7x=9.4 x=9.4/4.7 x=2

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

thank you